IODP Proceedings    Volume contents     Search


Operations summary

Transit from Shimizu (port)

Expedition 348 began on 13 September 2013 and ended on 29 January 2014. The Chikyu left the port of Shimizu en route for Site C0002. Typhoon Man-yi’s approach on 14 September sent the Chikyu into waiting on weather (WOW) status; the Chikyu remained in WOW status until 16 September and then returned to Site C0002, arriving on 17 September. Upon return to Site C0002, operations began with an ROV dive in Hole C0002F, and then preparation for coring tests with the new coring tool began.

Hole C0002M

Pressure tests of the riser joints were required, so the riser running equipment was rigged down to prepare for testing the SD-RCB assembly. Although designated the “small-diameter” coring tool, the SD-RCB core is actually larger than standard RCB cores (7.3 versus 6.6 cm inner diameter); the bit itself has a smaller diameter, 8½ inch versus the standard 10½ inch (or larger) drill bits. The coring BHA (Table T4) was run into the hole on 19 September 2013. Hole C0002M drilling began from 30.5 mbsf (1996.5 m BRT). By 21 September, the BHA had drilled ahead to 475 mbsf, and coring began from that depth (Tables T1, T2), cutting four cores and ending at 512.5 mbsf (final total depth). After coring operations were complete, the BHA was pulled out of hole and recovered to the surface on 22 September.

Hole C0002F

The rig up for riser running began, and the ROV dove to deploy transponders. Pressure tests of riser joints continued until all were secured for another standby at Shionomisaki for WOW until 25 September. The Chikyu returned to the Site C0002 blow-out preventer (BOP) running point, 15 mi northwest of Site C0002 and prepared to run the BOP and riser. From 26 September 2013, the BOP was run down while low-pressure (300 psi) tests of the auxiliary line (kill and choke lines) began. The tests failed and troubleshooting began; it was not until 28 September that all the lines passed low-pressure and high-pressure (5000 psi) tests. The BOP, however, was landed on the BOP cart for transit to Mikawa Bay on 29 September to pick up the repaired riser joint. The Chikyu returned to the Site C0002 BOP running point once loading was completed. On 1 October, however, the approach of Tropical Storm Fitow returned the Chikyu to WOW status until the storm track and an unrelated weather front were observed. The LWD tools were loaded from supply boat during WOW.

Full operations resumed on 5 October; however, two typhoons (Wipha and Francesco) passed by the site, requiring standby and standoff with the disconnected lower marine riser package (LMRP) from the wellhead. Operations resumed on 26 October, when the LMRP was run into the water. The vessel moved to the well center (water depth = 1967.5 m BRT, 1939 m mud depth below sea level [MSL] and successfully landed and locked the LMRP on the BOP by 31 October. BOP Yellow control pod (POD) pressure tests and Blue POD function tests were conducted on 1 November. The 17 inch LWD and GeoPilot BHA was made up and run in the hole on 2 November, drilling through the two cement plugs. Drilling continued to 850.5 mbsf, when a pressure test of the 20 inch casing was conducted.

Hole C0002N

After circulation for hole cleaning, the sidetrack for Hole C0002N began at 860.3 mbsf on 4 November 2013. Sidetracking continued until the drilling inclination and azimuth reached 5° and 0°, respectively. On 5 November at 1191.5 mbsf, inclination was changed to 0°. There were many occurrences of mud loss, overpull, and WOW while drilling both Holes C0002N and C0002P (Table T5). Drilling halted on 9 November (at ~1856 mbsf) when the real-time MWD signal was lost. After several attempts to recover the signal, it was decided to continue drilling without real-time monitoring to 2008.5 mbsf. On 10 November, the LWD BHA was pulled to the rig floor to remove the GeoPilot from the BHA, dump the LWD memory data, replace the MWD pulser, and perform maintenance on the mud pump. The new 17 inch LWD BHA was run into the hole, and drilling resumed from 2008.5 mbsf on 11 November. Target depth was reached at 2330 mbsf on 13 November. On 15 November the LWD BHA was pulled out of the hole, laid down, and LWD memory data were downloaded. Preparations to run and cement the 13⅜ inch casing began. On 18 November, the casing hanger joint was picked up and the 13⅜ inch casing was run in hole (RIH). Running 13⅜ inch casing continued to 1993.5 mbsf without observing significant drag. Tight spots were encountered while running the casing string, and the casing became stuck at 2024.5 mbsf. Despite all efforts, the casing needed to be pulled back and shortened by 25 casing joints, and a new pressure-assisted drill pipe running tool assembly was made up on 21 November. Running the shortened 13⅜ inch casing began; no drag larger than 100 kN was observed while running to 1502 mbsf. However, minor mud loss was observed at 1994.5 mbsf. The casing hanger was landed onto the wellhead on 22 November with the 13⅜ inch casing shoe set at 2010 mbsf. The casing was cemented, followed by a 5500 psi pressure test. The BOP Blue POD function and pressure tests were completed by 24 November.

The 12¼ inch drill-out cement BHA was made up, run into the hole, and tagged the top of cement at 1897.7 mbsf on 25 November. After a short halt in operations for WOW, operations resumed with 1.13 specific gravity (sg) drilling mud. Top of cement was tagged, mud was circulated bottoms up, and a casing pressure test began. The pressure test was completed, and preparations for drilling out cement started. On 27 November, drilling confirmed the presence of cement at 2008.5 mbsf, and rig up for the shoe bond test began. The test was successful, and the drill string was pressurized to an equivalent mud weight of 1.25 sg for 5 min. Drilling out cement continued, but when drilling reached 2028.4 mbsf, weight on bit (WOB) and the Hydralift power swivel (HPS; National Oilwell Varco) torque became unstable, fluctuating and suddenly increasing. The pipe was stuck, and attempts to free the pipe by jarring began, pausing only for derrick inspection. The Schlumberger free point indicator tool was rigged up to test for the exact stuck point in preparation for freeing the drill pipe by controlled explosion to allow the pipe to be backed off at a specified joint. The back-off tool broke the drill pipe at 1886.35 mbsf. The colliding tool was recovered, the drill string was pulled to the surface on 30 November, and the back-off point was confirmed and examined. Although seemingly undamaged, the pipe joint was marked and removed from use. Recovery attempts by fishing for the remaining portion of the BHA began from 1 December. After the fish had been jarred 724 times up and 173 times down with no signs of success, the Schlumberger free point indicator tool was rigged up again and confirmed the free pipe point at 1963 mbsf. The colliding tool was run into the hole to 1960.58 mbsf, calibrated the firing depth, and fired. Circulation and bottom-up circulation from 4 December was completed while the Schlumberger wireline tools were rigged down. The collided drill collar was recovered on deck on 5 December, and a quick examination of the drill collar pin end found that its outer diameter had expanded by inch (from 8½ to 9 inches); the pin thread also expanded and cracked.

The decision was made to kick off through the 13⅜ inch casing and begin a new sidetrack to continue drilling. The casing was scraped with a scraper BHA and then pulled out of the hole and laid for Schlumberger cement bond log wireline runs to confirm cement levels outside the casing, which is required for the window sidetrack milling. The cement bond log was run on 7 December and confirmed cement to 1352.5 mbsf. Time was spent with a milling BHA to certify the fish depth (1948 mbsf) before setting a bridge plug and whipstock, and beginning window cutting in the 13⅜ inch casing. Testing with a wireline run found a slight discrepancy (~2 m) in bridge plug depths, so another run with the casing collar locator wireline confirmed the top of plug depth at 1949.5 mbsf. Wireline recovery and rig down was completed early on 9 December.

Between 9 and 11 December, the Chikyu performed an emergency disconnect sequence and dismantled the HPS for a magnetic particle inspection of the main shaft of the main power unit. On 11 December, the Chikyu returned to the well to land the LMRP and return to milling out with the whipstock BHA. Once it was set at 1945.5 mbsf on 14 December, window milling in the 13⅜ inch casing began.

Hole C0002P

Hole C0002P begins at 1936.5 mbsf, the top of the milling window. Milling out continued to 1954.5 mbsf, stopping to ream and dress the window so that the BHA could freely pass through. After dressing, two LOTs were run. After completion, mud weight was increased to 1.18 sg. The milling BHA was pulled from the hole, and the kick-off BHA was made up and run into the hole on 16 December 2013. The 12¼ inch sidetrack kick-off BHA began alternate rotation and sliding forward to increase inclination and offset from Hole C0002N, while mud weight was raised to 1.23 sg. As with Hole C0002N, several short periods of stuck pipe and WOW occurred during drilling (Table T5). Mud weight was raised to 1.28 sg, as the kick-off inclination finally reached 3.9° by 2107.5 mbsf (sensor depth, 16.35 m above the bit) and dropped off by 2162.5 mbsf. After sweeping out the hole, the kick-off BHA was pulled to the rig floor on 21 December. RCB coring operations (Table T1) ran from 22 to 24 December (2163.0–2218.5 mbsf; 6 cores; Tables T2, T4). The RCB BHA was pulled from the hole and laid down on 25 December and replaced with a 12¼ inch LWD BHA that was run into the hole on 26 December. However, on 27 December, when the bit reached 2263 mbsf, MWD data telemetry stopped. All troubleshooting attempts failed, so from 28 December it was decided to drill ahead without real-time monitoring. Drilling (with short wiper trips) proceeded to 3058.5 mbsf by 31 December.

Hole C0002P was advanced with relatively little trouble during drilling to total depth, but had numerous overpulls and packoffs during wiper trips and hole cleaning operations (Table T5); each time, the pipe was freed after dropping pumping pressure and working the drill pipe. Drilling of the hole to 3058.5 mbsf was completed by 1 January 2014; a heavy load of cavings continued to arrive at the shale shakers even when not making new hole. The HPS stalled three times while working pipe while making connections, and mud weight was increased to 1.32 sg. On 2 January, the LWD BHA was recovered on deck and memory data were downloaded. Gamma ray, annular pressure-while-drilling, and resistivity data were recovered by 0545 h, and azimuthal sonic data were recovered by 0615 h, but the azimuthal focused resistivity data recovery port was damaged, preventing data download. Data download required shipping the azimuthal focused resistivity tool to the Halliburton base in Thailand.

Reaming out the open hole to 14½ inches was performed with the Anderreamer underreamer BHA from 2 January. No logging tools were part of this BHA. The underreamer began reaming open the 12¼ inch hole to 14½ inches from 1963.5 mbsf. Reaming down occasionally paused for mud pump maintenance. Hole opening continued, with nothing more than traces of gas and cavings encountered throughout this period of drilling. From 8 January, several episodes of HPS stalling at 2867.5 mbsf from heave increase were noted. Another cold front weather system required pulling out of the hole for WOW to above the BOP for standby. The cold front passed, and WOW continued until 9 January. On 10 January, reaming up and down encountered frequent HPS stalls, tight spots, and one hole pack-off at 2902.5 mbsf. The pack-off initially had no mud return, but continuously working the pipe and reducing pumping pressure eventually freed the pipe by 1500 h. Backreaming and hole opening operations continued to 12 January, when the Anderreamer underreamer BHA was recovered on deck; the reamer cutters were excessively worn and damaged. A new underreamer BHA was rigged up and run into the hole. Once the BHA reached 1957.5 mbsf, the Anderreamer was activated. A series of issues with the mud pumps impacting drilling, with frequent pauses in drilling to swab out and replace cylinders of the mud pumps. Reaming downhole resumed and by 14 January had reached 2960.5 mbsf. Hole opening had progressed 4 m when the hole packed off, leaving the drill pipe stuck in the hole. Operations to free the pipe began immediately, with firing jars, spotting lube, and pulling until the pipe was freed. Sweeping the hole began and circulation and bottoms-up was finished by 15 January. The BHA continued in a series of wiper trips and reaming, reaching 2964.5 mbsf by 2130 h. Circulation and bottoms-up to clean the hole began and continued until 16 January. It was decided to spot Hi-vis mud (30 m3) and pull the underreamer to the surface; the BHA was laid down by 1800 h.

With reaming and drilling complete, running and cementing the 11¾ inch liner began. On 18 January, the liner was run to 966.5 m MSL while setting the guide shoe joint, float collar joint, landing collar joint, and the 3½ inch ball in the float collar joint. No drag was observed while running casing. Slack was checked and weight picked up before circulating after breaking pipe to drop the 1¾ inch ball. Setting the liner hanger began; the first three attempts to set the hanger failed. After picking up 1 m and applying pressure, it was confirmed that the hanger had been released. However, it soon became clear that the liner-casing packer had also set when circulation attempts failed. Pressure tests confirmed that the annulus pressure was holding (up to 7.8 MPa), and there was no communication with the formation below. Two LOTs were performed, the second ending in an injection test. The liner packer was pressure tested at 5000 psi for 5 min and then at 4000 psi for 15 min; both tests passed. After the LOTs the liner hanger running tool was pulled to surface, laying out the cement stand on 19 January. The first of two cement “squeezes” was completed early on 20 January. After squeezing, the diverter was pulled from the hole and Hi-vis mud was spotted. The diverter assembly was pulled to surface, and the cement plug was pressure tested on 21 January.

Riser recovery began, and the BOP was disconnected from the wellhead on 22 January while the ROV set the corrosion cap and recovered transponders. By 23 January, the gooseneck and termination joints had been recovered and riser joints proper began to be recovered; the BOP was recovered and set on the BOP cart by 25 January. The ROV dove to complete transponder recovery. Once complete, the ROV was recovered to the surface while riser-handling tools were rigged down and general rig equipment services were performed. The Chikyu left the site on 29 January for facilities in Irago, Aichi Prefecture (Japan).